Fermi National Laboratory
 

Accelerator Update for January 19, 2007 - January 22, 2007

 

Friday January 19

The day shift began with Operations monitoring the Tevatron (TeV) store 5186 (69.41E30) and a Recycler stash (290.3E10), with the Antiproton Source (Pbar) stacking (18.4E10), with Linac, Booster, Main Injector (MI), and Recycler operating normally, with MiniBooNE off, with NuMI taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

At 8:31 AM, MI aborted for the first of seven times in about five hours due to bad beam positions. It reset each time.

A MI expert began a two-hour study at MI-60 at 10:52 AM.

At 11:26 AM, a MI RF station (MIRF13) tripped off. It reset okay.

At 12:25 PM, the NuMI horn tripped off on a ground fault indication. It reset okay.

Pbar experts began a stacking study at 1:09 PM.

Operations received the Meson MTest run conditions at 1:31 PM.

Operations reported at 4:06 PM that the Recycler had stashed twice during the day shift.

A TeV expert began an end-of-store study around 3:03 PM.

Operations terminated store 5186 at 4:14 PM.

Pbar experts finished their stacking study at 4:25 PM. Normal stacking resumed.

Operations put the TeV into shot setup at 4:31 PM.

At 5:26 PM, Operations began loading antiprotons into the TeV. An emittance blowup occurred on transfers 7 and 9. Operations called in an expert.

Operations established store 5187 at 6:05 PM with an initial luminosity of 235.5E30.

At 10:37 PM, a Recycler expert reported that he made an adjustment to the gain of the transverse dampers.

Operations reported at midnight that the Recycler had stashed twice during the evening shift.

Saturday January 20

The midnight shift began with Operations monitoring store 5187 (104.92E10) and the stash (113.9E10), with Pbar stacking (48.0E10), and with NuMI taking beam.

At 3:43 AM, a Linac RF station (LRF2) tripped off on a “driver crowbar” and wouldn’t reset. This trip halted beam. Operations called in an expert.

At 4:04 AM, MI suffered a MI-10 vacuum burst. Operators reset the vacuum pumps. It took about fifteen minutes to recover the vacuum and open the valves.

At 4:33 AM, MI-8 suffered a vacuum cascade. It took about ten minutes to recover.

At 6:44 AM, the Tevatron Electron Lens (TEL) tripped off on a vacuum interlock due to a pump (L1VAC1) tripping off. Everything reset okay.

Operations reported at 7:56 AM that the Recycler had stashed twice during the midnight shift.

At 10:56 AM, Operations reported that a Recycler emittance blowup had caused the loss of about 15mA of antiprotons. Operations called in experts.

At 1:48 PM, Linac experts reported that they were replacing the cathode for the 4616 screen modulator.

Operation reported that technicians had completed LRF2 work and that operations had established beam to the Linac absorber at 3:32 PM.

NuMI resumed taking beam at 3:34 PM.

At 5:09 PM, Operations reported that a NuMI Lambertson (I:LAM61) had tripped off on a rack filter summation indication. The supply wouldn’t reset. Operations contacted a NuMI expert and then called in a power supply expert.

A MI expert began a barrier bucket study at 5:51 PM.

Operations reported that the power supply expert had completed his work on the NuMI Lambertson power supply. NuMI resumed taking beam at 6:22 PM.

At 7:24 PM, Operations reported that LAM61 was tripping off due to timing input problems. Operators began replacing the McClure timing input box at 7:24 PM.

At 8:19 PM, Operations reported several more NuMI trips. Operators consulted with experts and tried various repairs. The trips stopped, but Operations couldn’t say that they had fixed the problem.

Stacking resumed at 8:52 PM.

NuMI resumed taking beam at 9 PM.

Operations terminated store 5187 at 9:08 PM and then put the TeV into shot setup.

At 9:39 PM, a Pbar Debuncher ENI status for DRF1-6 had tripped off and wouldn’t reset. Much later, operators investigated but couldn’t get the device to reset. Since stacking didn’t seem to be suffering much, Operations decided to let it go until a Pbar expert could look at it in the morning.

At 10:14 PM, Operations reported that some of the protons injected into the TeV from MI were going into the wrong buckets. Operations contacted an expert who instructed an operator to make an adjustment that improved coalescing. Operations completed loading the protons about 45 minutes later.

At 11:14 PM, Operations reported that they were only getting about 60% of the antiprotons out of the Recycler and into the TeV. Recycler experts weren’t sure of the cause. They advised Operations to continue the injection.

Operations reported at midnight that the Recycler had stashed once during the evening shift.

Sunday January 21

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash and stack, and with NuMI taking beam.

Operations established store 5188 at 12:05 AM with an initial luminosity of 64.77E30. Recycler experts reported that they would investigate the extraction problems in the morning.

The Run Coordinator supplied Operations at 12:35 AM with two damper plots. The first plot showed the bad emittances from store 5188. The second plot showed a good store from the past.

At 1:37 AM, the H- Source started having reference trips. Operators investigated and reported regulation problems. They reset the source locally.

At 1:51 AM, the H- Source tripped off for the first of many, many times over the rest of the midnight shift on an overcurrent indications. Operators locally reset the source each time.

A Recycler expert arrived at 7:19 AM to investigate the damper problems.

Operations reported at 8:02 AM that the Recycler had stashed twice during the midnight shift.

At 10:36 AM the H- Source went down and wouldn’t reset. Operations called in experts.

Operations reported at 11:42 AM that they were able to turn on DRF1-6.

A Recycler expert had begun damper studies earlier this morning. He completed his study around 2:17 PM. They reported that they didn’t find anything conclusive.

Operations reported around 4 PM that the Recycler had stashed once during the day shift.

TeV experts began a collimator end-of-store study around 5:01 PM.

Operations established beam to the Linac absorber at 6:04 PM.

NuMI resumed taking beam at 6:19 PM

Pbar resumed stacking at 6:23 PM.

Operations terminated store 5188 at 6:52 PM.

Operations put the TeV into shot setup at 7:05 PM.

At 7:55 PM, the TeV sector B4 separator reversing switch got stuck. Operators were able to reset it locally.

At 8:47 PM, Operations reported that the H- Source problems continued. Operators called an expert who said the column was sparking. He suggested turning the voltage up very slowly, which is what the operators did.

Operations put the TeV through a dry squeeze at 10:47 PM.

A Linac expert arrived at 11:03 PM to help with the H- Source.

Monday January 22

The midnight shift began with the TeV in a dry squeeze, and with no beam anywhere else due to H- Source problems.

At 12:34 AM, the Preaccelerator expert accessed the H- Source, but he didn’t find any obvious problem.

At 2:01 AM, the Preaccelerator expert got the source running with a voltage of -666 kV. This provided about 16mA of beam out of Linac.

Beam resumed to NuMI at 2:03 AM.

Operations put the TeV into shot setup at 2:12 AM.

The Preaccelerator expert reported that the H- Source was running at -665kV to minimize sparking. He adjusted other 750keV and Linac devices to run at the source’s lower voltage.

Operations established store 5190 at 3:49 AM with an initial luminosity of 111.89E30.

At 4:55 AM, a TeV sector B17 horizontal separator tripped off. Operators investigated.

Operations contacted TeV experts and made a plan for turning on the B17 separator. The experiments turned off their detectors. At 5:44 AM, Operators tried, but couldn’t turn on the B17 separator.

At 6:02 AM, operators reported that the separator’s power supply had failed.

Operations terminated store 5190 at 7:37 AM.

Operations reported at 8:06 AM that the Recycler had stashed twice during the midnight shift.

The Plans for Monday

The plans for today are to fix the B17 separator and do some other small repairs in the TeV tunnels. Later this morning Preaccelerator experts will replace the H- water resistor. The I- Source won’t be ready to use for at least several days. The Run Coordinator will allow approved maintenance work on the other machines.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 01/22/2007    email Fermilab